### v1

# Forecasting CPI inflation under economic policy and geo-political uncertainties

Abstract
Forecasting a key macroeconomic variable, consumer price index (CPI) inflation, for BRIC countries using economic policy uncertainty and geopolitical risk is a difficult proposition for policymakers at the central banks. This study proposes a novel filtered ensemble wavelet neural network (FEWNet) that can produce reliable long-term forecasts for CPI inflation. The proposal applies a maximum overlapping discrete wavelet transform to the CPI inflation series to obtain high-frequency and low-frequency signals. All the wavelet-transformed series and filtered exogenous variables are fed into downstream autoregressive neural networks to make the final ensemble forecast. Theoretically, we show that FEWNet reduces the empirical risk compared to single, fully connected neural networks. We also demonstrate that the rolling-window real-time forecasts obtained from the proposed algorithm are significantly more accurate than benchmark forecasting methods. Additionally, we use conformal prediction intervals to quantify the uncertainty associated with the forecasts generated by the proposed approach. The excellent performance of FEWNet can be attributed to its capacity to effectively capture non-linearities and long-range dependencies in the data through its adaptable architecture.
1 Introduction
The monetary policy framework at central banks relies heavily on forecasts of various macroeconomic variables. Forecasts have also grown in importance as a means of communication for central banks during the past 20 years, and they have the potential to shape public and market expectations. Numerous studies have examined central bank forecasting from various angles due to the paramount significance of these predictions[1, 2, 3, 4, 5]. Forecasting, particularly in the context of different macroeconomic policy variables, is one of the major focus areas in this domain. Inflation is one such policy variable considered instrumental in designing an economy’s overall monetary policy. Thus, a deeper understanding of the factors that help forecast inflation accurately is fundamental to policymakers at the central bank [1]. Over the past few decades, emerging market economies such as Brazil, Russia, India, and China (BRIC) have undergone significant changes in their macroeconomic conditions, including shifts in monetary policy frameworks such as inflation targeting, increased globalization, and the impact of events like the financial crisis and the recent pandemic. These shifts make predicting key macroeconomic variables like consumer price index (CPI) inflation quite challenging [2]. Given these dynamic changes in the macroeconomic environment, factors like economic policy uncertainty and geopolitical risk might offer critical insights to forecast inflation accurately.
Due to the significance of inflation forecasting exercise for monetary policy, central banks often depend on distinct models, namely structural and non-structural. The structural models, which are different versions of Phillips curve models, are derived from economic theory. On the other hand, the non-structural models, such as linear univariate and multivariate time-series models, aim to utilize the unique characteristics of the data without explicitly depending on any economic theory [3]. Non-structural models often prioritize forecast accuracy above causal inference and are particularly valuable for short-term, out-of-sample forecasting. Linear time series models such as autoregressive integrated moving average (ARIMA) and vector autoregressive (VAR) are limited in their ability to account for business cycles, periods of extreme volatility caused by macroeconomic uncertainties, and structural changes. As a result, non-linear, hybrid, and machine learning models have become increasingly popular to address these shortcomings. Furthermore, the existing linear time series forecasting methods bring several undesirable properties, ranging from high sensitivity to model specification to high-frequency data requirements [4]. This is especially important for forecasting inflation or similar macroeconomic variables that are only available at the monthly, quarterly, and annual frequency. This issue is more serious for emerging market economies where the quality of macroeconomic data is not up to the mark. Given such issues related to inflation data in emerging countries, machine learning (ML) models and combined approaches provide an opportunity to improve forecast performance.
Recently, an increasing number of studies have been conducted on applying ML models for inflation forecasting, both for advanced and emerging market economies. A study conducted in the United States utilized a basic feed-forward neural network to forecast quarterly CPI inflation. The study concluded that this approach was successful in predicting CPI inflation [5]. Neural networks proved more effective than traditional models in predicting the monthly inflation rate for the OECD countries [6]. Their research revealed that neural network models outperformed simple AR models in 45% of the nations, on average, whereas simple AR models performed better in 23% of the countries. Recurrent neural networks, in their various forms, have been discovered to be beneficial for predicting inflation in developed countries [7, 8, 9, 10]. In addition, a recent study compared various machine learning models, including lasso regression, random forests, and deep neural networks, to analyze inflation in the United States. The study also considered several external factors, such as cash and credit availability, online prices, housing prices, consumer data, exchange rates, and interest rates [11]. Their research has shown that machine learning models with many covariates consistently outperformed conventional techniques. This superiority can be attributed to nonlinear relationships between historical macroeconomic variables and inflation. This has dispelled doubts about using machine learning and contemporary deep learning architectures in analyzing big data in economics [12]. For emerging market economies, the study of inflation forecasting using advanced ML models is in a nascent stage. Applications of neural networks are evident in developing countries like India, Brazil [13], Russia [14], Turkey, South Africa [15], and China [16]. However, none of the above studies, to the best of our knowledge, have considered exogenous factors like macroeconomic uncertainties for inflation forecasting, especially for emerging markets.
Following events such as the global financial crisis, it has been observed that many countries worldwide have encountered heightened uncertainty. This has had an impact on the decision-making process of individuals and consequently on the overall macroeconomy [17, 18, 19]. Although previous studies have primarily examined the connection between uncertainty and the overall economy, the evaluation of the changing relationship between measures of uncertainty and inflation is still in its early stages. It has not yet yielded a definitive conclusion[20, 21]. Gaining insight into the intricacies of inflation dynamics and uncertainty metrics is quite critical for policymakers. Multiple recent studies indicate that both economic policy uncertainty and geopolitical risk significantly influence the movement of inflation within a given sample [22, 20, 23, 24, 25, 26]. Consequently, inflation forecasting models that do not incorporate these uncertainties are deemed to be incorrectly specified. The study conducted by [23] examines the effectiveness of different uncertainty measures in predicting inflation. In particular, they use US inflation data and the economic policy uncertainty index (EPU) to compare how well the vector autoregressive fractionally integrated moving average (VARFIMA) model performs at making predictions compared to other models. The study conducted by [27] examined the predictive impact of Economic Policy Uncertainty (EPU) on stock market volatility in the United States and the United Kingdom. The findings suggest that EPU has a significant positive effect on volatility. However, this positive relationship weakens when the signal quality is low. [28] asserted that Economic Policy Uncertainty (EPU) has a crucial impact on enhancing the predictive accuracy of models forecasting stock market volatility. The study conducted by [29] investigated the impact of policy uncertainty in the United States and Europe on the short-term pricing model for gold. The study conducted by [30] investigated the impact of Economic Policy Uncertainty (EPU) on predicting future real economic activities in the UK and European economies. For an extensive analysis and thorough examination of different methods for quantifying uncertainty, as well as the impact of economic policy uncertainty on various economic activities such as stock market returns, corporate capital investments and spending, corporate finance, and risk management, please refer to the publication by [31]. On the other hand, the effect of geopolitical risk uncertainty (GPRC) has been studied extensively by [24]. A recent study by [32] aims to link geopolitical risk with the reduction of firm values as well. The studies covering the impact of various uncertainty measures on different economic activities are still in the nascent phase, and this study in turn, presents a way to accommodate the roles of EPU and GPRC in forecasting inflation numbers.
Wavelet analysis has shown remarkable promise in analyzing financial time series, due to its ability to distillate crucial information from the financial data. It is their structural flexibility that allows them to handle irregular time series data more effectively for long-term time series forecasting [33]. Its ability to spot the time regime shifts and breakpoints allows it to discover important signals within the time series data. Broadly, wavelets are functions that can decompose the original time series into two sub-components: one concerning the low-frequency or deterministic component and the other concerning the high-frequency or stochastic component of the series. This process aims to reduce the effect of ‘noise’ in the forecast and is found to be quite useful in dealing with a variety of non-stationary signals, a common feature of financial and macroeconomic time series data, as they are constructed over finite time intervals and can represent the ‘local’ elements in both time and scale [34]. Several applications of wavelet analysis can be found in forecasting exchange rates, GDP growth, crude oil prices, stock market prices, Taiwan stock index, S&P dividend yield, trade prices, expenditure and income, price movements, money growth, and inflation [35, 36], volatility in foreign exchange markets, car sales from the last decades [37]. A wide variety of hybrid models combining wavelets and neural networks have been studied for forecasting natural gas prices [38], future trading [39], short-term electricity load [40, 41], energy price forecasting [42], oil prices [35], stock index [43]. Apart from forecasting, the application of wavelet analysis can be found in various areas of finance: [44] analyzed the association between stock price and inflation rate using wavelet analysis, [45] explored the aspect of assessment of the risk in the emerging markets, [46] studied the value at risk in the metal markets. [47] examined the impact of the global financial crisis on the multi-horizon nature of systematic risk and market risk using a wavelet multi-scale approach. [48] used wavelet neural networks for the prediction of corporate financial distress. However, none of the above studies have utilized inflation series and its causal counterparts for long-term forecasting.
Against this background, the primary aim of this study is to create a filtered ensemble wavelet neural network (FEWNet) model. This model belongs to a novel category of machine learning models that combine traditional neural networks with wavelet models. The objective is to generate accurate long-term forecasts for CPI inflation in BRIC countries while taking into account economic and geo-political uncertainties. FEWNet can produce reliable long-term forecasts for CPI inflation of BRIC countries conditioned on economic and geopolitical uncertainties. In analyzing macroeconomic policy variables like CPI inflation, the wavelet decomposition can extract the low-frequency components of the series without sacrificing the fundamental characteristics of the original time series. This model also integrates a layer for feature engineering aspects by extracting the trend components using the Hodrick-Prescott (HP) filter and cyclical elements using the Christiano-Fitzgerald (CF) ideal band-pass filter in the forecasting exercise. Initially, the observed time series (CPI inflation) and exogenous factors (EPU and GPRC) undergo decomposition into their trend and cyclical components using HP and CF filters in the first phase. Subsequently, in the second phase, all the trend and cyclical components are modeled as auxiliary inputs within the FEWNet framework, along with the wavelet-transformed details and smooth coefficients of the CPI inflation. This comprehensive approach aims to generate forecasts for 12 months and 24 months ahead. Conditioning inflation forecasts on policy uncertainty and risk uncertainty would capture the inflation or deflation effects of uncertainty. We argue using wavelet coherence analysis that there exists a relationship between EPU and GPRC with that of CPI inflation for BRIC economies. We show that the data-generating properties of inflation (persistence and volatility) are associated with the movements in EPU and GPRC. Proposed FEWNet generates out-of-sample forecasts in a recursive manner, and using several statistical metrics and tests, we compare its performance with alternative forecasting methods. Theoretical results on empirical risk minimization are also presented for the FEWNet. Apart from improvement in forecast accuracy measures of CPI inflation for the BRIC countries, another critical area that interests the policymakers is the knowledge of how various forms of structural uncertainties can influence the movement or complex dynamic behavior of different macroeconomic variables. The performance of FEWNet is superior to its counterparts in all three economies, except for China. This demonstrates that FEWNet is highly effective in handling complicated, non-stationary time series in forecasting tasks.
The paper is organized in the following way. Section 2 explains the macroeconomic series considered in this study, various data pre-processing strategies, and the causal association between the uncertainty indices and CPI inflation. Section 3 focuses on the proposed methodology with a detailed overview of its architectural design and Section 4 theoretically establishes the robustness of the method from an empirical risk minimization perspective. Section 5 discusses the experimental evaluation and the process of deriving a conformal prediction interval for the FEWNet framework. Finally, Section 6 ultimately wraps up the study by presenting a thorough analysis of the main discoveries and outlining potential future directions for the research.
2 Data and Preliminary Analysis
This work is premised on forecasting the monthly CPI inflation series for the four countries: Brazil, Russia, India, and China based on their historical data and various uncertainty indices EPU and GPRC [^2] Two news-based measures of uncertainty are used in this study: the geopolitical risk uncertainty index (GPRC) [24] and the economic policy uncertainty index (EPU) of [19]. EPU stands for Economic Policy Uncertainty, which represents the likelihood of changes in monetary, fiscal, or regulatory policies that can impact the decision-making behavior of investors and consumers [19]. The geopolitical risk index[24] measures the occurrence of geopolitical events such as terrorism, local and regional political instability, political violence, coups d’état, territorial conflicts, and war. These events contribute to the uncertainty in the geopolitical landscape. Please refer to http://www.policyuncertainty.com/ and https://www.matteoiacoviello.com/gpr.htm for more information on the EPU and GPRC index calculations.. To evaluate the performance of our proposed method, we consider two forecast horizons of length 12 months and 24 months for each of the countries under study. In this section, we discuss the macroeconomic data considered in our analysis and their characteristics, along with the data preparation step (Section 2.1) and the causality analyses (Section 2.2) performed in this study. Furthermore, the implementable codes and the datasets used in this study are made available on GitHub repository for the reproducibility of the results.
2.1 Data Preparation and Global Charateristics
[…]
$\pi_{\text{Jan }}^{2003}=\left[\frac{CPI_{Jan}^{2003}-CPI_{\text{Jan }}^{2002}}{CPI_{Jan}^{2002}}\right]*100\%,$
where $\pi_{Jan}^{2003}$ denotes the derived CPI inflation number for 2003-Jan, calculated using CPI values of 2003-Jan and 2002-Jan (prior year and same month). In addition to the CPI inflation, we also consider the impact of two uncertainty measures, namely the EPU and GPRC index, on the future projections of CPI inflation numbers for the BRIC countries. As part of the data pre-processing step, we perform the log transformation on the EPU series to stabilize its variance. Sequentially, we decompose the CPI inflation, log-transformed EPU, and GPRC series using the Hodrick and Prescott (HP) filter [49] and the Christiano-Fitzgerald (CF) random walk filter [50]. These economic filters play a vital role in extracting the trends and business cycles associated with the time series. Further details on the HP filter and CF filter are provided in Appendix 7.2.
In the analysis, we study the global features of the CPI inflation series, log-transformed EPU series, and GPRC series to identify their structural patterns and detect appropriate techniques for long-term forecasting. Primarily we focused on six popular time series characteristics, including skewness, kurtosis, non-linearity, long-range dependence, seasonality, and stationarity [51]. To assess the long-range dependency of the time series, we utilize the Hurst exponent. To analyze the non-linear characteristics of the series, we employ Tsay’s test and Keenan’s one-degree test for non-linearity. To determine the stationarity of the series, we conduct the Kwiatkowski–Phillips–Schmidt–Shin test. Lastly, to identify any seasonal patterns in the series, we apply the Ollech and Webel’s test. The results of the statistical features, summarized in Table 1, depict that most of the macroeconomic time series are non-stationary except for the GPRC series of Brazil and the CPI inflation data of China. The majority of these series are positively skewed except for the log-transformed EPU series of Russia. Furthermore, most of the economic time series exhibit non-linear patterns which plays a crucial role in selecting the appropriate modeling approach for the dataset. Additionally, based on Ollech and Webel’s seasonality test and the Hurst exponent we observe that all the series are non-seasonal and are characterized by long-range dependency. Moreover, the ACF and PACF plots of the CPI inflation series (presented in Table 2) depict the presence of serial autocorrelation. These inherent patterns of the CPI inflation series motivate the design and application of hybrid forecasting approaches for long-term forecasting.
Table 1: Global characteristics of the economic time series under study for BRIC countries.
Countries Series Skewness Kurtosis Non-Linearity Long-Range Dependence Seasonality Stationarity
Brazil CPI Inflation 1.67 3.57 Non-linear 0.73 Non-seasonal Non-stationary
log(EPU) -0.9 0.29 Non-linear 0.77 Non-seasonal Non-stationary
GPRC 2.06 6.60 Non-linear 0.63 Non-seasonal Stationary
Russia CPI Inflation 0.28 -0.98 Linear 0.78 Non-seasonal Non-stationary
log(EPU) -0.03 -0.12 Non-linear 0.79 Non-seasonal Non-stationary
GPRC 1.41 2.46 Linear 0.79 Non-seasonal Non-stationary
India CPI Inflation 0.71 0.17 Linear 0.82 Non-seasonal Non-stationary
log(EPU) 0.12 -0.47 Linear 0.80 Non-seasonal Non-stationary
GPRC 2.51 15.22 Non-linear 0.73 Non-seasonal Non-stationary
China CPI Inflation 0.82 1.08 Non-linear 0.69 Non Seasonal Stationary
log(EPU) 0.23 -0.87 Non-linear 0.82 Non-seasonal Non-stationary
GPRC 1.28 1.51 Non-linear 0.80 Non-seasonal Non-stationary
Table 2: Training data for FEWNet including target series CPI inflation (blue) with corresponding ACF, PACF plots and exogenous variables log-transformed EPU (orange) and GPRC (grey) for the BRIC countries.
Country
Training data
ACF plot
PACF plot
Brazil [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Russia [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
India [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
China [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2.2 Causality Analysis
This section explores the causal impact of log-transformed EPU and GPRC uncertainties on the CPI inflation series for the BRIC countries. Several statistical methods, such as the Granger causality test, cross-convergent mappings, and transfer entropy, have been suggested in the literature to examine the causal relationship between two variables [52]. Nevertheless, the tests rely on certain data-level assumptions that are not met in the macroeconomic series analyzed in this work. In order to address this constraint, we employ wavelet coherence analysis (WCA) to examine the causal relationship between the series [53]. The WCA methodology offers a valuable method for examining the interconnection and simultaneous movement of two non-stationary signals in the time-frequency domain.
Figure 1: Wavelet coherence analysis plots - CPI inflation, log-transformed EPU (top) and CPI inflation, GPRC (bottom) for BRIC Countries
In the WCA plots, demonstrated in Figure 1, the first row depicts the WCA results between CPI inflation and log-transformed EPU series, and the second row shows the same for CPI inflation and GPRC series. In each of the plots horizontal axis measures the time, and the vertical axis shows the frequency dimension. Wavelet coherence represents the regions in the time-frequency domain where the two signals co-vary. In the plot, the transition from warmer colors (red) to colder colors (blue) indicates regions from higher coherence to lower coherence between the series. The arrow in the WCA plots indicates the lag phase relations between the analyzed signals. A zero-phase difference would hint toward the co-movement of the signals on any particular scale. The direction of an arrow has a specific connotation, with the arrows pointing to the right (left) indicating that the signals are in phase (anti-phase), i.e., the two signals are moving in the same direction or vice versa. As evident from the plots, there exists a significant causal relation between CPI inflation, log-transformed EPU, and GPRC series. This in turn, justifies the choice of EPU and GPRC as exogenous factors in forecasting the CPI inflation series of BRIC countries.
3 Proposed FEWNet Model
Figure 2: MODWT decomposition of the CPI Inflation series for (a) Brazil, (b) Russia, (c) India, and (d) China between the period Jan-2003 and Nov-2021. In these figures, $\delta_{1},\delta_{2},\ldots,\delta_{5}$ and $S_{5}$ represent the details and smooth of the series generated by the MODWT-based MRA approach. The bottom chart in all the figures represents the original time series (CPI inflation) in actual frequency scale
This section provides a comprehensive description of our proposed architecture, called the filtered ensemble wavelet neural network (FEWNet). FEWNet combines the maximal overlapping discrete wavelet transformation (MODWT) technique (explained in Appendix 7.1), economic filtering methods such as the Hodrick-Prescott (HP) filter and Christiano-Fitzgerald (CF) filter (explained in Appendix 7.2), and the autoregressive neural network with exogenous variables (ARNNx). The FEWNet framework is a sequential approach that initially decomposes the CPI inflation data into low-frequency (smooth) and $\mathcal{K}$ high-frequency (detail) components using the scaling and wavelet filters of MODWT, as depicted in Figure 2. The ability of the transition and time-invariant MODWT approach to handle the non-stationary features of the time-dependent signal makes it a valuable tool in pre-processing the CPI inflation dataset. Simultaneously, to capture the economic uncertainties of the time series, we apply the CF and HP filters on the CPI, log-transformed EPU, and GPRC series. The non-parametric CF filter efficiently captures the trend components, whereas the HP filter is suitable for identifying and modeling the business cycles. This long-term trend and the repetitive cyclical patterns have been found to provide significant information regarding the economic uncertainties that impact the future trajectories of the CPI inflation series.
The real-world Consumer Price Index (CPI) inflation series exhibit notable features, including non-stationarity, nonlinearity, and long-term dependence, as evidenced in Table 1. In order to accurately represent this non-uniform time series, we utilize MODWT decomposition using Daubechies orthogonal wavelets of length 2, namely the Haar filter. This transformation has the ability to identify the specific time and frequency components of signals, successfully distinguishing between long-range connections, non-stationarity, and patterns within the signals [54]. By utilizing the MODWT-based multi-resolution analysis (MRA) decomposition on the series $Y_{t}$, we derive a collection of $\mathcal{K}+1$ distinct random variables called wavelet and scaling coefficients, which are not correlated with each other. In our suggested methodology, we employ separate neural networks within an ensemble framework to represent $\mathcal{K}+1$ random variables, instead of employing the original CPI inflation series $Y_{t}$. These neural networks are specifically engineered and trained to effectively manage the intricacies linked to the wavelet coefficients. By employing this approach, we may efficiently tackle the non-linear nature observed in the Consumer Price Index (CPI) inflation data of the BRIC nations. In addition, in order to capture the temporal causal association between the input series and the exogenous variables ($X_{j,t},\;j=1,2,\ldots,6$), we include them in each of the $\mathcal{K}+1$ wavelet transformed series of $Y_{t}$ in their corresponding neural networks. The forecasts produced by these $\mathcal{K}+1$ neural networks are combined to determine the future dynamics of $Y_{t}$. The mathematical expression for the $m$-step ahead forecasts of $Y_{t}$, designated as $\hat{Y}_{\mathcal{N}+m}$, based on $\mathcal{N}$ past data, is as follows:
$\hat{Y}_{\mathcal{N}+m}=\hat{S}_{\mathcal{K},\mathcal{N}+m}+\sum_{\tilde{k}=1}^{\mathcal{K}}\hat{\delta}_{\tilde{k},\mathcal{N}+m},$ (1)
where $\hat{S}_{\mathcal{K},\mathcal{N}+m}$ and $\hat{\delta}_{\tilde{k},\mathcal{N}+m},\;\tilde{k}=1,2,\ldots,\mathcal{K}$ are the forecasts generated from the simultaneously computed neural networks of the smooth series $S_{\mathcal{K},t}$ and the details series $\delta_{\tilde{k},t}\;\tilde{k}=1,2,\ldots,\mathcal{K}$. Each of these $\mathcal{K}+1$ neural networks consists of a feedforward design with three layers - the input layer with $p$ nodes, one hidden layer with $q$ nodes, and one output layer without any shortcut connections. The neural networks model incorporates $p-6$ lagged observations of the wavelet coefficients, as well as one lagged observation for each of the six exogenous variables ($X_{j}$). These variables represent the trend and cycle information of CPI, log-transformed EPU, and GPRC. This setting allows the networks to generate one-step forward forecasts for the related series. Mathematically, the result obtained from each neural network after one iteration can be represented as:
$\begin{gathered}\hat{\delta}_{\tilde{k},\mathcal{N}+1}=\alpha_{0,\tilde{k}}+\sum_{i=1}^{q}\beta_{i,\tilde{k}}\sigma(\alpha_{i,\tilde{k}}+\beta^{{}^{\prime}}_{i,\tilde{k}}\vec{\delta}_{\tilde{k}}+\gamma^{{}^{\prime}}_{i,\tilde{k}}\underline{X});\;\tilde{k}=1,2,\ldots,\mathcal{K}\\ \hat{S}_{\mathcal{K},\mathcal{N}+1}=\alpha_{0}^{S}+\sum_{i=1}^{q}\beta_{i}^{S}\sigma(\alpha_{i}^{S}+{\beta^{S}_{i}}^{{}^{\prime}}\vec{S}_{\mathcal{K}}+{\gamma^{S}_{i}}^{{}^{\prime}}\underline{X}),\end{gathered}$
where $\vec{\delta}_{\tilde{k}}$ and $\vec{S}_{\mathcal{K}}$ denotes the $p-6$ lagged observation of the corresponding wavelet coefficients and $\underline{X}$ indicates a single lagged vector of exogenous variables, $\alpha_{0,\tilde{k}},\alpha_{0}^{S},\alpha_{i,\tilde{k}},\alpha_{i}^{S}$ are the bias terms, $\beta_{i,\tilde{k}},\beta_{i}^{S}$ are the connection weights between hidden and output layer, $\gamma^{{}^{\prime}}_{i,\tilde{k}},{\gamma^{S}_{i}}^{{}^{\prime}}$ are the weight vectors between the exogenous variables and the hidden layer, $\beta^{{}^{\prime}}_{i,\tilde{k}},{\beta^{S}_{i}}^{{}^{\prime}}$ are the connection weights between lagged wavelet series and the hidden layer, and $\sigma$ indicates the nonlinear activation function. Within our framework, we commence by assigning random initial values to the connection weights. Subsequently, we employ the gradient descent back-propagation approach to train these weights, as described in the work by [55]. The aforementioned technique produces a forecast that is one step ahead for each of the wavelets and scaling coefficients. In order to make forecasts for several future time steps, we incorporate the most recent predictions into the input layer instead of using the last observed value, and we continue this procedure recursively. Ultimately, we combine the predictions produced by each of the base models in an ensemble setup and derive the forecast for a for a desired time horizon.
[…]
where $\mathcal{V}$ indicates the validation set and $Y_{t}$ and $\hat{Y}_{t}$ are the original series and forecast at time $t$. Moreover, to stabilize the learning rate of the neural network, restrict overfitting, and to reduce the run-time of the proposed FEWNet framework, we set the number of hidden nodes arranged in a single hidden layer as $q=\left[\frac{p+1}{2}\right]$ following [51, 57]. The suggested FEWNet model is illustrated in Figure 3 using a schematic approach. The target time series is initially subjected to a MODWT-based MRA, resulting in multiple degrees of transformation, as shown in the figure. The economic uncertainties and geopolitical risk index series are filtered along with the CPI series using the CF filter and HP filter to generate the corresponding cycle (green circles) and trend (yellow circles) respectively. These six exogenous variables are then provided with the lagged wavelet series in the input layer of each of the $\mathcal{K}+1$ neural networks. The connection weights establish the connections between each input set and the hidden nodes, represented as lines with arrows. The results from the hidden layer are processed in the output layer, ultimately producing the forecast for the next step from the interconnected network. Ultimately, the future dynamics of the CPI inflation series are formed by producing multi-step ahead forecasts in a recursive fashion and then combining them in an ensemble. The utilization of wavelet decomposition to reveal the complex variability of the transformed series, and the subsequent modeling of these fluctuations with diverse economic filtered uncertainties using an auto-regressive neural network with auxiliary information within an ensemble framework, provides a valid rationale for naming the proposed model as FEWNet. The implementation methodology of the proposed FEWNet is outlined in Algorithm 1.
Figure 3: Illustration of the proposed FEWNet model
Input: A target time series $\{Y_{t},\;t=1,2,\ldots\mathcal{N}\}$ indicating CPI inflation and exogenous series $X_{t}$ comprising of log-transformed EPU, and GPRC.
Output: $m\;(\geq 1)$ step-ahead forecast of the CPI inflation series $Y_{t}$.
Methodology:
1 Transform the target data vector ($Y_{t}$) into $\mathcal{K}$ levels by applying MODWT with haar filter to generate a sequence of $\mathcal{K}+1$ uncorrelated random variables indicating the local trend and high-frequency fluctuations of the series.
2 Determine the value of $\mathcal{K}$ as $\operatorname{log}_{e}(\mathcal{N})$, where $\mathcal{N}$ represents the size of the training dataset.
3 To quantify the economic uncertainty parameters of inflation data, apply CF and HP filters on the exogenous variables and the target variable to generate their corresponding cyclical and trend behavior, respectively.
for $\tilde{k}\leftarrow 1$ to $\mathcal{K}$ do
       Utilize an auto-regressive neural network to model the $\tilde{k}^{th}$ detail series with $X_{j,t},\;j=1,2,\ldots,6$.
       Build a feed-forward neural network consisting of an input layer with $p$ nodes, a hidden layer with $q$ neurons, and an output layer.
       Input one-lagged observation for each of the six filtered covariates along with $p-6$ lagged values of the $\tilde{k}^{th}$ detail series as input to the neural network.The neural network will then pass this input via $q$ hidden neurons and generate the forecast for the next time step.
       Determine the value of $p$ that minimizes the SMAPE and set $q=\left[\frac{\displaystyle{p+1}}{\displaystyle{2}}\right]$, for stable learning architecture and shorter run time of the model.
       return One-step forward forecast of the details series.
4Apply an auto-regressive neural network to both the smooth series of $Y_{t}$ and the exogenous variables, following the same definition used for the details series. Use this model to generate a forecast for future values.
5 Create an ensemble framework to merge the forecasts produced from several series and get the final one-step forward prediction.
6 Iteratively generate multi-step forward forecasts.
Algorithm 1 Proposed FEWNet model
4 Empirical Risk Minimization
FEWNet utilizes wavelet decomposition as a filtering stage and combines it with autoregressive neural networks that incorporate exogenous variables. This combination allows for precise predictions of time series that are both non-stationary and nonlinear. Wavelet decomposition yields a hierarchical structure of new time series derived from the original time series, facilitating their modeling and forecasting. This section demonstrates that FEWNet effectively minimizes the empirical risk in comparison to classical methods, hence highlighting its practical relevance.
Let $Y_{t},\;(t=1,2,\ldots,\mathcal{N})$ be the observed time series (stationary if random) and we aim to forecast the values of $Y_{p+1},Y_{p+2},\ldots,Y_{p+m}$ ($m\geq 1$ is the forecast step) using $p$ observations iteratively. In regression setup, a functional relationship is established between $Y_{p+m}$ and the input vector [$Y_{1},Y_{2},\ldots,Y_{p}$] by minimizing the predictive error. The optimal prediction sequence $\hat{Y}_{p+1},\hat{Y}_{p+2},\ldots$ minimizes the generalization risk defined as follows
$\mathcal{R}_{\text{Gen}}=\lim_{T\to\infty}\frac{1}{T}\sum_{p=1}^{T}\mathbb{E}\left[\left(\hat{Y}_{p+m}-Y_{p+m}\right)^{2}|Y_{p},Y_{p-1},\ldots\right],$
where $\hat{Y}_{p+m}=\mathbb{E}\left[Y_{p+m}|Y_{p},Y_{p-1},\ldots\right]$. However, computation of $\hat{Y}_{p+m}$ is difficult as the observed time series has $\mathcal{N}$ observations (training samples). Therefore, we minimize the following empirical risk:
$\mathcal{R}_{\text{Emp}}=\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\left(Y_{i}-\hat{Y}_{i}\right)^{2},$ (2)
where $\hat{Y}_{i}$ is the prediction (expectations can be omitted for deterministic time series). In practical problems of economic time series forecasting, the relationship between $\hat{Y}_{p+m}$ and the observed time sequence $Y_{p},Y_{p-1},\ldots$ is supposed to be nonlinear in nature. Thus, the inclusion of autoregressive neural networks (a set of feed-forward artificial neural networks designed specifically for predicting future values in time series data) for forecasting is justified in this exercise. The relationship can be estimated by
$\hat{Y}_{p+m}=f(Y_{p},Y_{p-1},\ldots,Y_{1}),$ (3)
where $f$ denotes an autoregressive neural network function. The model’s order ($p$) is an essential parameter to estimate. Smaller values of $p$ make the model training simple but miss the information about the past lagged values whereas larger values of $p$ compel the model training hard resulting in a “curse of dimensionality”. To avoid the issues of underfitting and overfitting, we use the ARNNx model, where we choose $p$ by minimizing the SMAPE for the validation set. If $f$ in Eq. (3) is a complex estimator (e.g., deep learning models such as transformers, deep MLP, etc), it will learn well on the training data but perform poorly on unseen test data. Application of MODWT decomposes the time series using low and high pass band filters iteratively resulting in a trend series and several detailed series (containing dynamics of the economic system at different scales). ARNNx is then trained to model the trend and detailed series and provide predictions for future values. The combined values of these MODWT-decomposed series provide a prediction of the original economic data. Although the main motivation for using wavelet decomposition on these economic time series is to enhance the predictive power, MODWT gives the smooth term $S_{\mathcal{K}}$ that contains the slowest dynamics (noise-free) and $\delta_{\tilde{k}}$ that contains system dynamics (noisy). The low-resolution data are frequently contaminated by noise. Consequently, training an ARNNx model on these time series is less challenging compared to the original time series. As $\tilde{k}$ increases, the dynamics get slower, resulting in low-level detail series that consist solely of noise. To mitigate overfitting, one can assign a value of zero to their predictions.
Several recent works [56, 58, 59, 57] have proposed to treat wavelet-decomposed time series independently. Another complicated approach will be to treat each series along with other series (treated as exogenous variables) but it results in increasing the dimensionality of the time series forecasting problem. In the current formulation we model $S_{\mathcal{K}}$ and other $\delta_{\tilde{k}},\;(\tilde{k}=1,2,\ldots,\mathcal{K})$ series using ARNNx models as follows:
$\hat{S}_{\mathcal{K},p+m}=\hat{f}_{0}\left(S_{\mathcal{K},p},S_{\mathcal{K},p-1},\ldots,S_{\mathcal{K},p-r_{0}},\underline{X}\right)\text{ and }\hat{\delta}_{\tilde{k},p+m}=\hat{f}_{\tilde{k}}\left(\delta_{\tilde{k},p},\delta_{\tilde{k},p-1},\ldots,\delta_{\tilde{k},p-r_{\tilde{k}}},\underline{X}\right),\;\tilde{k}=1,2,\ldots,\mathcal{K}.$ (4)
The choice of $\hat{f}_{i}(i=0,1,2,\ldots,\mathcal{K})$ is related to the properties of the time series (nonlinear for this case). [57] proposed nonlinear ARNNx model and each $\hat{f}_{i}$ may have the same order for $r_{i}$ (we will use $r$ for the rest) to decrease the model’s complexity. The predictions are made as an equal-weighted combination of smooth and details predictions:
$\hat{Y}_{p+m}=\hat{S}_{\mathcal{K},p+m}+\sum_{\tilde{k}=1}^{\mathcal{K}}\hat{\delta}_{\tilde{k},p+m}.$
In this context, we establish empirical risk minimization (ERM) property for our proposed FEWNet framework. In statistical learning theory, ERM defines a family of learning models and provides theoretical bounds on their performances. This is useful since in practice we can’t generalize how well a model will work (called true risk) due to not knowing the true data distribution. However, we can instead measure its performance on a known training data set (called empirical risk).
In FEWNet, we use ARNNx estimators $\hat{f}_{0},\hat{f}_{1},\ldots,\hat{f}_{\mathcal{K}}$ with same order $r$ having equal number of neurons in the network architecture. These estimators are obtained by simultaneously minimizing the following empirical risk:
$\mathcal{R}_{\text{Emp}}^{W}=\frac{1}{\mathcal{N}}\sum_{p=1}^{\mathcal{N}}\left[\left(S_{\mathcal{K},p+m}-\hat{S}_{\mathcal{K},p+m}\right)+\sum_{\tilde{k}=1}^{\mathcal{K}}\left(\delta_{\tilde{k},p+m}-\hat{\delta}_{\tilde{k},p+m}\right)\right]^{2}.$ (5)
Since ARNNx estimators can be expressed as a nonlinear combination of linear projections of lagged inputs for the CPI inflation series, the following result holds:
Proposition 1
[…]
Proof 1
Using Eq. (3) we write, the ARNNx estimator as follows
$\hat{f}=\alpha+\sum_{j^{{}^{\prime}}=1}^{q^{*}}\beta_{j^{{}^{\prime}}}\sigma\left(z_{j^{{}^{\prime}}}\right),$
where $z_{j^{{}^{\prime}}}=\alpha^{*}+\sum_{i=1}^{r}\beta_{i,j^{{}^{\prime}}}^{*}y_{i}+\gamma^{*}\underline{X}$ where the parameters are estimated by minimizing Eq. (2). For FEWNet we have
$\begin{gathered}\hat{f}_{0}=\alpha_{0}^{S}+\sum_{i=1}^{q}\beta_{i}^{S}\sigma(\alpha_{i}^{S}+\sum_{j_{0}=1}^{r}{\beta^{S}_{i,j_{0}}}S_{\mathcal{K},j_{0}}+{\gamma^{S}_{i}}^{{}^{\prime}}\underline{X})\\ \hat{f}_{\tilde{k}}=\alpha_{0,\tilde{k}}+\sum_{i=1}^{q}\beta_{i,\tilde{k}}\sigma(\alpha_{i,\tilde{k}}+\sum_{j_{\tilde{k}}=1}^{r}\beta_{i,j_{\tilde{k}}}{\delta}_{\tilde{k},j_{\tilde{k}}}+\gamma^{{}^{\prime}}_{i,\tilde{k}}\underline{X});\;\tilde{k}=1,2,\ldots,\mathcal{K}.\end{gathered}$
[…]
Since $\hat{f}$ is a linear combination of $\hat{f}_{0}$ and $\hat{f}_{\tilde{k}},\;(\tilde{k}=1,2,\ldots,\mathcal{K})$ with uniform weights and scaling ($\tilde{g}_{l}$) and wavelet ($\tilde{h}_{l}$) filter satisfies unit energy assumption and even length scaling assumption in MODWT i.e.,
$\tilde{h}_{l}=-\tilde{g}_{l}\;\forall l\geq 1,\;\;\tilde{h}_{0}+\tilde{g}_{0}=1,\;\;\tilde{h}_{l}=\tilde{g}_{l}=0\;\forall l\leq-1$ (6)
Similar results may hold for any other decomposition method satisfying the Eq. (6) in fact.
Remark 1
However, it by no means guarantees the generalization risk reduction. In practical scenarios, the selection of the models for $\delta_{\tilde{k},p}$ and $S_{\mathcal{K},p}$ would be crucial and can be done using cross-validation. Also the choice of $r_{i},\;(i=0,1,\ldots,\mathcal{K})$ may change the results in practice as can be thought of as a future advancement of this study. Our theoretical results show the robustness of the wavelet decomposed approaches in our proposal from an ERM perspective.
5 Experimental Evaluation and Analysis of Results
We assessed the efficacy of the proposed FEWNet framework by comparing its performance to other baseline predictions derived from statistical, machine learning, and deep learning methodologies. In our experimental evaluation, we utilize a 5-fold time series cross-validation approach to train all the forecasting frameworks. We then generate forecasts for two different time horizons, specifically 12 months and 24 months, to demonstrate the generalizability of the setup. In this section, we briefly explain the baseline models (Section 5.1), evaluation metrics (Section 5.2), benchmark comparison, and experimental results (Section 5.3).
5.1 Baseline Models
We compared the FEWNet model and various baseline models, as well as a range of machine learning and deep learning algorithms that allow for the incorporation of exogenous factors. The evaluation incorporates the following competitive models:
  • Auto-regressive Integrated Moving Average with exogenous variable (ARIMAx) model is widely used in time series forecasting [60]. The ARIMAx $(p,d,q)$ framework captures the linear patterns of the time series by considering $p$ previous values of the target series, $q$ prior forecast errors, and historical values of the exogenous variable. This framework employs a differencing technique of order $d$ to guarantee the stationarity of the data. The coefficients of the linear ARIMAx model are estimated by minimizing the Akaike information criterion (AIC) value.
  • Seasonal ARIMAx (SARIMAx) is an extension of the ARIMAx model for seasonal time series datasets. The model has a similar architecture as the ARIMAx model with additional parameters for modeling the linear trend in the seasonal components of the series. Thus the SARIMAx $(p,d,q)(P,D,Q,s)$ framework models the non-seasonal components of time series with parameters $(p,d,q)$ and for the seasonal components of length $s$ it utilizes the parameters $(P,D,Q,s)$.
  • Auto-regressive Fractionally Integrated Moving Average with exogenous variable (ARFIMAx) generalizes the classical ARIMAx model by allowing the value of the differencing parameter $d$ to be a fraction i.e. $d\in(-0.5,0.5)$. ARFIMAx (p, d, q) process has been widely studied to model and forecast time series exhibiting long-range dependence. As evident in Table 1, the CPI Inflation series along with the economic uncertainty indices exhibit long-range dependency, thus the choice of using the ARFIMAx model seems reasonable for our analysis to account for slowly decaying auto-correlations.
  • Deep-learning based Auto-regressive (DeepAR) model is an advanced deep learning system designed specifically for forecasting time series data. This framework employs a recurrent neural network architecture to forecast future trajectories in a dataset that varies over time [61].
  • Neural Basis Expansion Analysis for Time Series (NBeats) is a dense neural network structure explicitly developed for time series prediction problems. This model consists of many blocks, each composed of two main layers. The initial layer is responsible for modeling the time series to replicate previous observations and produce forecasts. Conversely, the second layer focuses on re-modeling the discrepancies between the actual and predicted values generated by the first layer and adjusting the forecasted figures. In our testing, we set the number of blocks as 2 to decrease the computational complexity of the framework.
  • Auto-regressive Distributed Lag The autoregressive distributed lag (ARDL) model has been extensively employed in the field of econometrics to estimate the long-run and short-run dynamics of time series data [62]. The distributed lag components of the ARDL model are capable of simulating both the stationary and non-stationary time series components.
  • Lasso Regression is a regularized linear regression model that handles overfitting using the L1 penalty. The penalty term involved in the loss function of the LASSO model helps in shrinking the coefficients of the regressor variables with limited predictive power, thus leading to feature selection.
  • Xtreme Gradient Boosting (XGBoost) is a supervised machine learning algorithm that intends to predict the target variable through the ensemble of a set of weak learners [63]. These tree-based models are trained in a sequential fashion, where the model attempts to minimize the remaining error components of all the prior trees. During prediction, the sum of all the predicted values of the weak learners is reported. In our experimental analysis, we convert the forecasting task into a supervised prediction problem where the lagged inputs and the exogenous variables are the features and the subsequent forecast is the required label for the instance.
  • Temporal Fusion Transformer (TFT) is a deep learning time series forecasting model that leverages the multi-head attention layer to capture the complex temporal dynamics of multiple time sequences. One of the key benefits of the TFT framework is its ability to model and forecast both univariate and multivariate time series simultaneously. Furthermore, the architecture is capable of incorporating additional information in the form of time–variant and static exogenous regressors. In our experimentation, we set the number of attention heads as 3.
  • Wavelet ARIMAx (WARIMAx) is a modified version of the ARIMAx model that use the MODWT decomposition technique for the purpose of predicting time series data [59]. The model breaks down the signal that changes over time into several wavelet and smooth coefficients. It then uses an ARIMAx model to make predictions based on each of these series.The multi-step forecasts are derived by aggregating these candidate forecasts for a desired horizon.
  • Auto-regressive Neural Network with exogenous variable (ARNNx) is a generalization of the classical feed-forward neural network model for the auto-regressive time series process [64]. This model utilizes previous observations of the target time series and the exogenous variables as input to the network. The ARNNx model comprises a single hidden layer positioned between the input and output layers. The ARNNx $(p,k)$ model uses $p$-lagged input values from the input layers and passes them via $k$ neurons in the hidden layer. The value of $k$ is typically calculated by the formula $k=\left[\frac{p+1}{2}\right]$, where $p$ is a given parameter. This approach helps mitigate overfitting and maintain a stable learning architecture [51]. The model gets initialized with random values and trained using the gradient descent back-propagation approach [55].
5.2 Evaluation Metrics
In our study, we have considered five evaluation metrics to assess the performance of the proposed forecaster and the baseline models in generating the 12-month and 24-month ahead forecasts of the CPI inflation series for the BRIC countries. These metrics include the Root Mean Square Error (RMSE), Mean Absolute Scaled Error (MASE), Symmetric Mean Absolute Percentage Error (SMAPE), Theils’ U1 measure, Median Relative Absolute Error (MDRAE), and Median Absolute Percentage Error (MDAPE). The mathematical expression of these metrics is provided as follows:
[…]
where $y_{t}$ is the ground truth, $\hat{y}_{t}$ is the forecast of the corresponding model, $\hat{y}_{t}^{R}$ is the naive forecast at time $t$, and $m$ is the forecast horizon. By convention, the model with the lowest value of the error metric is considered as the best-performing model [51, 57].
5.3 Experimental results and baseline comparison
This section focuses on the implementation and effectiveness of the proposed FEWNet architecture in forecasting the Consumer Price Index (CPI) inflation series of the BRIC countries. In order to conduct a thorough assessment, we compare the performance of FEWNet to various cutting-edge algorithms. The FEWNet model is implemented using the R statistical software. The MODWT technique is first implemented using the modwt function from the ‘wavelets’ package. This function decomposes the training data into wavelet and scaling coefficients using the pyramid algorithm with the ‘haar’ filter. The determination of the number of decomposition levels is done conventionally as the floor function of $\log_{e}$(length(train)) [34]. In the subsequent phase, each wavelet (details) and scaling (smooth) coefficient series is modeled with trend and cycle information from the log-transformed EPU, GPRC, and CPI data obtained through economic filters. This is achieved using an ARNNx model fitted with the nnetar function from the ‘forecast’ package in R. To select the number of inputs ($p$) for the individual networks, a time series cross-validation approach is applied with a search range of 1-24. Subsequently, these $p$ inputs are processed through $k=(p+1)/2$ hidden nodes arranged in a single hidden layer. This process generates one-step ahead forecasts for the individual series, which are aggregated to produce the final forecast. The framework iteratively generates multi-step ahead forecasts. The hyperparameters of the FEWNet($p,k$) algorithm used in this study are summarized in Table 3. Furthermore, for implementing classical forecasters such as ARIMAx, SARIMAx, ARDL, ARFIMAx, and WARIMAx, the ‘statsmodels’ library in Python is employed. Deep learning models, including DeepAR, NBeats, and TFT, are implemented using the ‘darts’ library in Python. The machine learning frameworks like Lasso and XGBoost are implemented using the ‘skforecast’ package in Python.
Country Algorithm $(p,k)_{FH=12M}$ $(p,k)_{FH=24M}$
Brazil FEWNet (6,5) (18,11)
Russia FEWNet (1,2) (18,11)
India FEWNet (18,11) (12,8)
China FEWNet (18,11) (12,8)
Table 3: Model parameters: $(p,k)$ of FEWNet for BRIC countries across different forecast – horizons. The table shows optimal parameter combinations for 12-month and 24-month rolling window forecasts of CPI inflation.
After implementing the proposed FEWNet model and other baseline architectures, we generate forecasts for the specified time horizons. Tables 4 and 5 display the performance of the models for the semi-long term (12-month) and long-term (24-month) timeframes, respectively, using out-of-sample data. The error metrics reported in the tables indicate that, in the majority of forecasting tasks, the proposed FEWNet framework outperforms the baseline models. In the semi-long term forecasting of Brazil, the FEWNet model significantly reduces the RMSE metric of the ARNNx model. This improvement in model performance is attributed to the application of wavelet decomposition in the FEWNet model. A similar conclusion is supported by other accuracy metrics. For the CPI inflation series of Russia, while the DeepAR model provides the best forecasts among the baseline models, the FEWNet model’s forecasts closely align with the ground truth. The 12-month ahead CPI inflation forecasts of India generated by the proposed architecture show substantial improvement over the best-performing NBeats and WARIMAx models. When forecasting the CPI series of China, the NBeats model competes with the FEWNet architecture, with the former exhibiting the best accuracy for all metrics except SMAPE. The Chinese Consumer Price Index (CPI) series demonstrates a pattern of stable trends (trend-stationary). While the wavelet-decomposition approach is commendable for tackling structural volatility and non-stationarity, its efficacy may be restricted when applied to trend-stationary or covariance stationary series. Consequently, it may not outperform other statistical or machine-learning methods when used on inherently stationary series (a property that is not prevalent for most macroeconomic variables). This explains the cause of the somewhat competitive or less efficient performance of FEWNet in predicting Chinese inflation numbers. In long-term forecasting, the FEWNet architecture consistently outperforms baseline models for all countries except China, where the NBeats model demonstrates more accurate forecasts of the CPI series. Throughout the experimental evaluations, the importance of wavelet transformation and the economic filtered auxiliary information becomes evident in efficiently capturing both time and frequency-level dynamics of the time series. While machine learning and deep learning models generally produce competitive performances for semi-long forecast horizons, their effectiveness diminishes significantly over the long-term horizon. In contrast, the FEWNet model consistently performs well over both horizons, demonstrating its generalizability. Additionally, the wavelet decomposition of the FEWNet approach proves to be a suitable technique for modeling time-frequency information, especially for target variables exhibiting non-stationary behavior (unlike China). Given that most economic time series follow a non-stationary trajectory, the FEWNet approach emerges as an ideal framework for their long-term forecasting even in the presence of structural volatility.
Table 4: Performance of the proposed FEWNet model in comparison to baseline forecasting techniques for 12 months ahead forecasts with exogenous factors EPU, and GPRC (best results are made bold).
Country Metrics DeepAR ARNNx NBeats ARFIMAx SARIMAx ARIMAx LR ARDL XGBoost TFT WARIMAx Proposed
FEWNet
Brazil RMSE 3.41 2.23 2.95 4.67 2.37 2.69 2.19 3.26 2.40 5.41 3.13 1.30
MASE 5.24 3.32 4.41 7.49 3.77 4.29 3.45 5.57 3.62 8.51 4.88 1.80
SMAPE(%) 37 25 34 68 29 33 26 52 27 85 39 12
Theil’s $U_{1}$ 0.20 0.14 0.22 0.38 0.16 0.19 0.14 0.25 0.17 0.48 0.23 0.09
MDRAE 4.42 2.45 5.22 8.38 3.84 4.44 3.70 5.24 4.41 9.26 5.45 1.97
MDAPE 0.30 0.20 0.34 0.59 0.27 0.30 0.23 0.40 0.26 0.46 0.39 0.10
Russia RMSE 2.00 4.16 2.49 12.68 6.65 7.80 4.45 3.65 6.21 3.87 3.72 0.88
MASE 4.66 10.17 5.63 33.84 17.62 20.26 11.28 9.67 16.06 9.96 9.44 2.21
SMAPE(%) 26 46 38 100 69 76 48 44 63 81 44 13
Theil’s $U_{1}$ 0.15 0.25 0.23 0.50 0.34 0.38 0.26 0.22 0.33 0.42 0.23 0.07
MDRAE 4.23 10.20 5.00 40.54 18.31 23.76 12.70 12.14 15.49 7.57 9.15 2.88
MDAPE 0.33 0.51 0.32 1.97 1.05 1.22 0.71 0.57 0.99 0.59 0.56 0.11
India RMSE 3.55 1.69 1.09 1.86 2.27 2.31 2.46 1.88 1.69 1.67 1.19 0.88
MASE 6.61 2.68 1.82 2.92 3.72 3.88 4.19 2.96 3.07 2.35 2.00 1.41
SMAPE(%) 52 33 22 28 51 54 58 38 38 24 23 15
Theil’s $U_{1}$ 0.27 0.21 0.12 0.17 0.29 0.30 0.33 0.23 0.21 0.16 0.13 0.09
MDRAE 9.18 2.18 1.70 2.43 3.91 5.16 4.26 2.37 3.93 1.86 2.46 1.22
MDAPE 0.68 0.26 0.20 0.38 0.30 0.34 0.44 0.30 0.36 0.15 0.22 0.15
China RMSE 1.33 2.01 0.90 10.45 4.54 5.21 2.35 3.34 0.99 2.11 4.85 1.70
MASE 1.73 2.76 1.18 15.68 6.68 7.67 3.24 4.75 1.25 2.77 7.18 1.93
SMAPE(%) 94 111 116 171 146 150 115 132 99 153 149 85
Theil’s $U_{1}$ 0.42 0.49 0.34 0.83 0.68 0.71 0.53 0.62 0.35 0.80 0.70 0.43
MDRAE 1.29 2.60 1.28 15.94 7.27 8.04 3.41 5.17 1.25 3.95 7.38 1.75
MDAPE 1.66 1.70 0.87 14.68 6.80 7.67 2.26 5.07 0.93 1.99 7.21 0.99
Table 5: Performance of the proposed FEWNet model in comparison to baseline forecasting techniques for 24 months ahead forecasts with exogenous factors EPU, and GPRC (best results are made bold).
Country Metrics DeepAR ARNNx NBeats ARFIMAx SARIMAx ARIMAx LR ARDL XGBoost TFT WARIMAx Proposed
FEWNet
Brazil RMSE 6.16 3.56 3.11 4.27 4.13 4.49 3.67 2.86 2.56 3.58 4.11 2.31
MASE 11.45 6.68 5.10 6.04 7.67 8.27 6.93 5.23 3.62 6.11 5.93 4.12
SMAPE(%) 91 57 47 58 64 69 58 47 40 55 57 39
Theil’s $U_{1}$ 0.45 0.27 0.29 0.47 0.32 0.36 0.27 0.23 0.21 0.31 0.45 0.19
MDRAE 12.66 6.81 5.60 3.67 9.31 8.31 7.49 5.87 5.11 6.85 3.76 4.84
MDAPE 0.86 0.60 0.48 0.36 0.56 0.65 0.70 0.44 0.41 0.50 0.39 0.36
Russia RMSE 3.26 6.12 5.26 11.69 9.90 9.49 6.97 6.67 2.16 2.72 5.22 2.24
MASE 8.39 18.32 14.63 35.12 28.78 27.25 20.31 19.75 5.10 7.00 15.03 5.28
SMAPE(%) 51 81 66 112 101 98 84 84 36 44 72 27
Theil’s $U_{1}$ 0.28 0.38 0.34 0.54 0.51 0.50 0.42 0.40 0.25 0.24 0.35 0.18
MDRAE 8.29 21.94 15.12 41.21 30.43 26.85 24.53 24.47 4.76 7.21 17.53 3.51
MDAPE 0.44 1.33 1.09 2.49 2.02 1.87 1.38 1.47 0.35 0.43 1.14 0.38
India RMSE 1.86 2.22 4.18 2.72 2.16 2.02 2.36 2.53 1.95 6.17 2.35 1.04
MASE 2.68 2.90 5.83 4.20 3.02 2.83 3.61 2.98 2.80 9.48 3.39 1.49
SMAPE(%) 26 35 46 38 34 31 47 37 33 65 37 17
Theil’s $U_{1}$ 0.15 0.23 0.29 0.21 0.21 0.19 0.26 0.26 0.20 0.37 0.21 0.10
MDRAE 3.25 4.31 6.22 6.46 4.02 4.07 4.21 2.12 4.28 13.47 3.70 1.42
MDAPE 0.31 0.24 0.49 0.42 0.30 0.28 0.41 0.21 0.26 1.07 0.40 0.14
China RMSE 1.71 1.85 1.43 18.37 3.48 4.04 1.59 2.00 2.87 2.05 7.40 2.01
MASE 1.95 2.15 1.75 25.72 4.23 5.05 1.89 2.33 3.12 2.42 10.17 2.37
SMAPE(%) 81 77 77 168 102 109 75 79 115 100 139 82
Theil’s $U_{1}$ 0.40 0.34 0.27 0.81 0.48 0.51 0.30 0.34 0.65 0.46 0.64 0.36
MDRAE 1.53 2.32 1.60 26.02 4.43 4.99 2.09 2.51 2.68 2.32 10.88 2.50
MDAPE 0.57 0.85 0.63 11.56 2.05 2.41 0.55 0.99 1.14 0.84 4.84 0.84
5.4 Statistical significance of the results
In this section, we assess the effectiveness of different forecasting models in terms of the statistical significance of measurement error differences using the model-agnostic multiple comparisons with the best (MCB) procedure [65]. This non-parametric test computes the average rank for each of the $\mathcal{M}$ forecasters based on their ex-ante accuracy across $\mathcal{D}$ datasets and identifies the model with the minimum average rank as the ‘best’ performing technique. The MCB test determines critical distances (CD) for each algorithm as $\Theta_{\alpha}\sqrt{\frac{\mathcal{M}\left(\mathcal{M}+1\right)}{6\mathcal{D}}}$, where $\Theta_{\alpha}$ is the critical value of the Tukey distribution at level $\alpha$, and it considers the CD of the best-performing model as the reference value for the test. Following the MCB test, we calculate the average rank of the algorithm and present the results of the MCB test for the RMSE metric in Figure 4. The plot reveals that the FEWNet framework has achieved the lowest mean rank of 2.12 among all competing algorithms for both forecast horizons. Therefore, it is deemed to be the best-performing model, followed by NBeats (4.38), XGBoost (4.50), ARNNx (5.25), and so forth. Furthermore, the upper limit of the critical distance for the FEWNet model (shown by the shaded area in the graphs) acts as the reference value for the test. When comparing SARIMAx, ARIMAx, and ARFIMAx to the baseline, it is observed that their critical intervals are much higher than the reference value and do not coincide with FEWNet. This indicates that their forecast performance is notably worse than the FEWNet approach. Our analysis underscores the statistical significance of the performance difference and the superiority of the proposed FEWNet model across datasets and forecast horizons. Moreover, the test establishes that FEWNet consistently and significantly outperforms other baseline models, as well as various machine learning or deep learning models.
Figure 4: Visualization of the multiple comparisons with best (MCB) analysis for (a) Brazil, (b) Russia, (c) India (d) China [BRIC] countries: In the figure, for example in (a) FEWNet- 2.12 means that the average rank of the proposed algorithm FEWNet based on RMSE error metric is 2.12, the same explanation applies to other countries and algorithms
5.5 Conformalized Prediction Intervals
Next, we focus on quantifying the uncertainty associated with the forecast generated by the FEWNet approach using a conformal prediction interval. The non-parametric conformal prediction approach, introduced in [66], is a set of methods that usually take an uncertainty score and turn it into a probabilistic band, including the true outcome. In other words, this framework converts point estimations into a prediction region. One of the significant advantages of conformal prediction intervals is that this model-agnostic procedure guarantees coverage. In the time series setup, this method generates the prediction interval by leveraging the sequential ordering of the data. Given the training set $\left\{Y_{t},\tilde{X}_{t}\right\}_{t=1}^{N}$ where $Y_{t}$ is the target series, and $\tilde{X}_{t}$ is the set of features, including lagged historical values of $Y_{t}$ and the covariates, we fit the FEWNet framework and an uncertainty model $\widehat{\Xi}$ on $\tilde{X}_{t}$ to generate a scaler notion of uncertainty. The conformal score $\mathcal{S}_{t}$ can be calculated as:
$\mathcal{S}_{t}=\frac{\left|Y_{t}-\operatorname{FEWNet}\left(\tilde{X}_{t}\right)\right|}{\widehat{\Xi}(\tilde{X}_{t})}$
Owing to the sequential nature of the data points, the conformal quantile for time series data can be calculated as a weighted conformal technique with a fixed $\kappa$-sized window $\omega_{t^{{}^{\prime}}}=\mathbb{1}\{t^{{}^{\prime}}\geq t-\ \kappa),\forall\ t^{{}^{\prime}}<t$. This now yields the quantile values as,
^$\widehat{\mathbb{Q}}_{t}=inf\left\{q:\ \frac{1}{\min\left(\kappa,t^{{}^{\prime}}-1\right)+1}\sum_{t^{{}^{\prime}}-1}^{t-1}{\mathcal{S}_{t^{{}^{\prime}}}\mathbb{1}\left(t^{{}^{\prime}}\geq t-\ \kappa\right)\geq 1-\alpha}\right\}$
Based on these weights-adjusted quantiles, the prediction interval at each time step $t$ can be computed as,
$\mathbb{C}\left(\tilde{X}_{t}\right)=\left[\operatorname{FEWNet}\left(\tilde{X}_{t}\right)-{\widehat{\mathbb{Q}}}_{t}\widehat{\Xi}\left(\tilde{X}_{t}\right),\operatorname{FEWNet}\left(\tilde{X}_{t}\right)+{\widehat{\mathbb{Q}}}_{t}\ \widehat{\Xi}\left(\tilde{X}_{t}\right)\right].$
In this section, we present the conformal prediction intervals for the 24-month ahead CPI inflation forecasts generated by the FEWNet model using the validation data points (2019-12 to 2021-11). The conformal prediction interval of the FEWNet framework, calculated based on the forecast estimates and residuals for the validation set, is demonstrated in Figure 5 along with the point estimate of the forecast produced by FEWNet and the best-performing baseline models for the BRIC countries. As observed in the figure the prediction intervals vary in width across different datasets. For instance, in the case of the 24-month forecast for Brazil, the width of the interval averages around 8.18, for Russia, it hovers around 7.93, for India, it reaches the minimum of 3.88; and for China, it averages around 7.60. A similar analysis of the conformal prediction intervals for the semi-long term forecast horizon is provided in Appendix 7.3. The overall analysis thus demonstrates a form of uncertainty quantification for the forecast of CPI inflation series for BRIC countries and shows the structural shift in the measurement of prediction uncertainty across different countries.
Figure 5: The plot shows the ground truth CPI inflation data (red dots), forecast generated by the best (yellow) and second-best (orange) baseline models for each country, conformal prediction interval (green shaded), and forecast (blue) generated by the FEWNet model for the long-term forecast horizon in BRIC Countries
6 Conclusion and Discussion
Historically, emerging economies have exhibited higher and more enduring inflation rates when compared to developed countries. Most emerging economies have implemented an inflation-targeting monetary policy framework in which inflation serves as the intermediate target. This paper outlines a novel machine learning framework, namely FEWNet, that combines the MODWT process with the ARNNx model for long-term CPI inflation forecasting for the BRIC countries under macroeconomic and geopolitical uncertainties. The proposed method demonstrates consistent performances for 3 out of 4 major BRIC economies for both the semi-long run and long-run forecast horizons. Since the CPI inflation data of China presents a stationary behavior, our proposal failed to outperform benchmark forecasters that work exceptionally well on stationary time series. Moreover, for Brazil, Russia, and India it produces the most accurate performance, outperforming the classical and SOTA machine learning or deep learning algorithms. We also present a framework for time series feature engineering using different economic filters (HP and CF band-pass filters) that can be very easily leveraged in the proposed architecture, which supports the inclusion of exogenous regressors within its flexible design. A combination of the multiresolution analysis and autoregressive neural networks are used to predict nonlinearity, nonstationarity, and non-gaussian time series of inflation. MODWT decomposition produces new time series that contains system and slow dynamics that easily separate signal from noise. Each time series is further modeled with ARNNx. Theoretically, we show that under certain conditions, the proposed FEWNet reduces the empirical risk, resulting in improved forecast accuracy.
Although FEWNet is designed to predict CPI inflation of BRIC countries, it can be seamlessly extended for developed countries and can include other uncertainty variables like Twitter-based uncertainties, financial stress indicators, world uncertainty index, etc. This framework can also be leveraged for forecasting other macroeconomic variables like GDP, gold price, unemployment rate, and others. The FEWNet framework is more effective for generating forecasts for non-stationary and non-linear time series data that are evident in macroeconomics; therefore, it can potentially serve as a viable alternative to forecasting strategy for central banks. In light of our results, further analysis of the FEWNet’s performance in predicting other macroeconomic time series as well as inflation in a multivariate forecasting setting seems to be a promising avenue for future research.